OpenStudy (anonymous):

Find the simplified form of the expression in scientific expression: (5 x 10^6)(6 x 10^-4)

OpenStudy (anonymous):

Can you help? @KamiBug

jhonyy9 (jhonyy9):

do you know the properties of exponents ?

OpenStudy (retireed):

Multiply coefficients 5 times 6 and exponents 6 to -4 .... What do you get?

jhonyy9 (jhonyy9):

x^a *x^b = x^(a+b) x^(-1) = 1/x for example

jhonyy9 (jhonyy9):

@retirEEd are you sure multiplie 6 by -4 will be right ?

OpenStudy (retireed):

Sorry I forgot the add thank you

jhonyy9 (jhonyy9):

ATTENTION please !!!

OpenStudy (retireed):

Multiply coefficients 5 times 6 and add exponents 6 to -4 .... What do you get?